# Prioritizing Security Practice Adoption: Empirical Insights on Software Security Outcomes in the npm Ecosystem

### v1

@@ 7. Discussion
The section discusses the novelty of our study (section 7.1) and the implications of our findings (section 7.2) and explores the threat to the validity (section 7.3) of our study.
7.1. Novelity of our study
Prior studies (zahan2023software, ; sonatype_8, ) have utilized OpenSSF Scorecard metrics alongside vulnerability count data to understand the relationship between security practices and security outcome metrics. Building on this foundation, we expand security outcome metrics to include MTTR and MTTU. Our findings showed that Vul_Count, MTTR, and MTTU are associated with security practices. However, MTTR is more complex and may have an influence on external factors such as project maintainability, resources, and dependency complexity.
Additionally, our study further advances existing research by addressing methodological limitations in prior work. Zahan et al. (zahan2023software, ) constructed a small dataset of 2,422 PyPI and npm packages and analyzed the relationship between OpenSSF Scorecard metrics and vulnerability counts. The authors’ findings revealed a positive association between security practice adoption and vulnerability counts, alongside a low explanatory power ($R^{2}$ 12%). The authors attributed these results to dataset size limitations, lack of vulnerable data, and the absence of control variables related to project characteristics. In contrast, our study leverages a larger and more representative dataset, encompassing approximately 146,951 npm package repositories, which mitigates biases introduced by small sample sizes and enhances generalizability. Additionally, we incorporate control variables, including project age, size, contributors, and dependency count and downloads, to account for potential confounding factors influencing security outcomes. Furthermore, Zahan et al.’s dataset had a temporal mismatch between security practice adoption and vulnerability data. The authors collected security practice scores at a single point in time but included all historically reported vulnerabilities throughout a package’s lifetime, potentially distorting the actual relationship between security practices and security outcomes. Our methodology addresses this inconsistency by aligning security practice adoption with vulnerabilities that actively existed at the time of our study. By ensuring a concurrent snapshot of security practices and vulnerabilities, our approach provides the temporally aligned evaluation of security practice effectiveness.
Additionally, we addressed the issue of data consistency in study (zahan2023software, ), stemming from repositories containing multiple packages, each potentially contributing different Vul_Count  (Section  3.2), despite sharing identical Scorecard scores. Moreover, vulnerabilities in prior studies (zahan2023software, ) did not include vulnerabilities from dependencies or entire repositories but rather only individual packages, introducing further inconsistencies. Such discrepancies result in duplicate predictors with varying outcome values, potentially biasing predictive models. To mitigate these issues, we filtered our dataset to include only repositories linked uniquely to a single package. We considered vulnerabilities explicitly associated with the repository itself and those inherited from dependencies.
In another study, Sonatype (sonatype_8, ) employed a random forest classification model to classify repositories as vulnerable or non-vulnerable using Security Scorecard metrics and project attributes. However, a binary classification approach is inherently limited in capturing relationships between security practices and security outcomes. The classification does not effectively measure the magnitude or direction of associations, nor does it indicate the relative strength or complexity of individual predictors. In contrast, our study utilizes regression and causal analysis, methods specifically designed to explore and quantify associations, thus providing insights into how variations in security practices systematically relate to security outcomes.
Summary: Overall, our methodological refinements, including extensive dataset size, the inclusion of project characteristics, causal analysis, targeted dataset inclusion strategies, and the introduction of MTTR and MTTU as additional novel security outcome metrics, offer a more precise and contextually generalizable evaluation of the relationship between security practices and outcomes.
7.2. Implications of our study
Our study provides actionable insights for practitioners, policymakers, and researchers by establishing measurable security outcome metrics to assess the effectiveness of security practices. For practitioners, our findings highlight that Code Review, Maintained, Pinned Dependencies, and Branch Protection are strongly associated with security outcomes. Organizations can prioritize these practices to improve their project security. By periodically tracking security outcome metrics, practitioners can evaluate whether their security interventions are effectively reducing vulnerabilities and improving remediation efficiency. Our research supports policymakers by providing empirical evidence on the effectiveness of security practices, informing data-driven policy decisions, and enhancing regulatory frameworks. For researchers, our study underscores the importance of incorporating longitudinal data, expanded covariate sets, and more granular temporal analyses to understand better the dynamic interactions between security practices, external factors, and security outcomes. Additionally, future research should explore new security outcome metrics beyond vulnerability counts and remediation times to develop a more comprehensive framework for measuring security improvements in software development.
7.3. Threats to validity
@@
Our regression and causal analyses provide evidence supporting associations between aggregated Scorecard scores, security practices, and security outcomes within our dataset. However, several inherent biases in our dataset must be acknowledged. First, our analyses are based on snapshot data capturing repository Scorecard metrics score and characteristics at a single point in time, lacking temporal granularity. Without longitudinal or time-series analysis, uncertainty remains about whether changes in security practices precede or directly influence observed security outcomes, thereby limiting generalizability beyond our dataset. Second, selection bias may be present, as repositories that adopt higher security practices could systematically differ from those with lower adoption in unmeasured aspects such as organizational resources, developer expertise, or security culture. The unobserved factors may confound the relationships we examine, leading to potential over or underestimations of the true impact of security practices. Third, although PSM attempts to mitigate confounding factors by simulating randomized experiments, PSM inherently relies on observed covariates. Therefore, unobserved or missed variables such as organizational policies, economic incentives, developer motivations, or project complexity could influence both security practices and security outcomes. Additionally, the consistently observed positive ATE for MTTR suggests that security practices may introduce operational overhead, and the relationship is context-dependent, where repositories with project characteristics may experience varying impacts.
@@
Apart from that, our dataset exhibits other imbalances, as most repositories have security practice scores of either 0 or -1, reflecting the limited adoption of these practices across the npm ecosystem. The real-world imbalance, characterized by few repositories adopting security practices, may introduce bias into our models, potentially skewing predictions or causal estimations. However, such imbalance accurately mirrors the current state of practice adoption within the npm, highlighting a widespread lack of security practice adoption among npm repositories. Hence, the imbalance reflects the actual state of adoption rather than a methodological flaw, highlighting the need for targeted interventions to increase security practice adoption among npm repositories. Finally, while our filtering approach improves data consistency by retaining only repositories uniquely linked to a single npm package and aligning vulnerability data at the repository level and MTTR and MTTU data at the package level, some inconsistencies may persist. For example, repositories that maintain multiple packages from different ecosystems or use different languages may still introduce variations in MTTR and MTTU value, potentially affecting model accuracy.
Summary: Overall, while our study provides insights into the relationships between security practices and security outcomes, the limitations highlight the need for further research. Future studies incorporating longitudinal data, expanded covariate sets, and more granular temporal analyses will be essential to better capture the dynamic interactions between security practices, external factors, and security outcomes.
8. Related Works
@@
This section highlights prior related studies, existing security guidelines, frameworks, and security outcome metrics. Security Practice Frameworks: Executive Order (EO) 14028 emphasized adopting and self-attesting security practices to improve software security. To align with the EO, government agencies and industry experts have introduced various software supply chain security frameworks to enhance software supply chain security. These include the Software Component Verification Standard (SCVS) by OWASP (scvs, ), the Building Security In Maturity Model (BSIMM) (BSIMM, ), the CNCF Technical Advisory Group (TAG) recommendations (CNCF_TAG, ), Supply Chain Levels for Software Artifacts (SLSA) (SLSA, ), NIST Secure Software Development Framework (SSDF) (souppaya2022secure, ), the Open Source Software (OSS) Secure Supply Chain (SSC) Framework by Microsoft (Microsoft_framework, ), NIST’s Cybersecurity Supply Chain Risk Management Practices (NIST800-161, ), the CISA Secure Software Development Attestation Form (CISASecureSoftwareAttestation, ) and OpenSSF Scorecard (Scorecard, ). The Proactive Software Supply Chain Risk Management (P-SSCRM) (Williams2024PSSCRM, ) released in 2024, results from the study of real-world software supply chain risk management initiatives and the union of the 73 security practices in ten government and industry documents (standards and frameworks) mentioned above. In 2025, EO 14144 (EO2025Cybersecurity, ) was issued, emphasizing the need for NIST to evaluate common cyber practices and security outcomes across industry sectors.
@@
Security Outcome Metrics: Various metrics can assess the effectiveness of security practices by tracking security performance over time, generally categorized as threat-oriented, asset/impact-oriented, and vulnerability-oriented (NIST800-55, ; NIST800-30, ). While threat- and asset/impact-oriented metrics offer valuable insights, they are difficult to quantify at scale due to reliance on real-time threat intelligence and complex impact assessments. In contrast, vulnerability-oriented metrics are more accessible and relevant to software security. Among them, severity-based metrics (e.g., CVSS) face challenges such as scoring inconsistencies and reliance on exploitability data (pendleton2016survey, ). Temporal vulnerability metrics, such as vulnerability count, time to detect, time to repair and time to remediate, are widely used to evaluate security (morrison2014mapping, ; sec_outcome, ; CISA-2023, ; morrison2018mapping, ). Morrison et al. (morrison2018mapping, ) conducted a systematic mapping study on software security metrics to evaluate the security properties of software and found that the most cited and most used metric is the vulnerability count (vulnerability metric). Davidson (davidson2009good, ) studied the time to close bug/vulnerability’ (a form of mean time to repair) as a security metric and measured its effectiveness. Beres et al. (beres2009using, ) studied the window of exposure’ (vulnerability metric), which is the length of interval of time needed by the security team to remediate a vulnerability once it has been disclosed. However, the scarcity of reported vulnerabilities, compounded by silently fixed issues, limits their reliability in assessing security outcomes (alhazmi2007measuring, ; dunlap2023finding, ). To address these limitations, prior work suggests alternative metrics, such as fault prediction models (shin2013can, ). Additionally, in software reliability, MTTR and MTTU have long been used to measure system resilience and maintenance efficiency (gokhale1999time, ; calvo2023applying, ). While MTTR is often employed in security contexts (MTTR_1, ; MTTR_2, ; MTTR_3, ), its reliance on reported vulnerabilities restricts its applicability. MTTU, by focusing on dependency updates, offers broader coverage but remains constrained by proprietary measurement methods (cogo2019empirical, ).
@@ 9. Conclusion
In this study, we investigated how aggregate Scorecard score and 12 Scorecard security practices metrics impact security outcome metrics. Our regression model and causal analysis reveal that higher aggregated Scorecard scores are associated with fewer vulnerabilities and shorter MTTU. However, while regression models indicate shorter MTTR, causal analysis suggests that factors beyond security practice adoption, such as project maintainability, resource availability, and dependency complexity, influence MTTR. Feature ranking highlights Code Review, Maintained status, Pinned Dependencies, and Branch Protection as top security practices associated with security outcomes, though their associations with security outcomes vary in direction. Additionally, MTTU exhibits trends similar to vulnerability count and MTTR, making it a potential proxy metric when vulnerability and MTTR data are unavailable. Our research aligns with federal cybersecurity initiatives (EO 14144), by providing empirical evidence to guide security policy and industry practices. Future work should incorporate longitudinal data and additional security outcome metrics to refine further our understanding of how security practices impact software security over time.

### v2

@@ 7. Discussion
The section discusses the research implications (section 7.1) and explores the threat to the validity (section 7.2) of our study.
7.1. Research Implication
Defining and validating security outcome metrics is challenging; our work represents a step toward advancing that goal. Our study provides actionable insights for practitioners, policymakers, and researchers by investigating measurable security outcome metrics to assess the effectiveness of security practices.
Holistic security practice adoption:. RQ1 demonstrates that aggregated scores consistently improve security outcomes, while RQ2 reveals context-dependent and sometimes counterintuitive effects of individual practices. Our findings suggest that comprehensive security practice adoptions are more effective than isolated measures, as coordinated practices generate synergistic effects that individual actions alone cannot achieve.
Context-Aware Adoption: The impact of security practice on security outcomes varies by project characteristics, requiring tailored strategies. While selecting dependencies, organizations should prioritize mature, actively maintained large projects with large contributors, fewer dependencies, and high downloaded packages, as our analysis showed that projects have fewer vulnerabilities and faster MTTR due to better security performance. Conversely, smaller or inactive packages with fewer downloads require enhanced monitoring and proactive security assessments to compensate for slower MTTU.
Strategic Practice Prioritization: Despite individual variability, Code Review, CI-Tests, License, Contributors, and Pinned Dependencies consistently demonstrate high feature importance across security outcomes. Our findings suggest that practitioners can adopt these empirically validated practices within their workflow, while recognizing that practices like Code Review, CI-test positive association reflect enhanced detection capabilities rather than security degradation. To ensure that security practice adoption is yielding tangible benefits, teams should regularly monitor outcome metrics. Continuous tracking of these indicators enables practitioners to assess the effectiveness of implemented practices and iteratively refine their security strategies based on empirical feedback.
Evidence-Based Policy Development: The quantified relationships between security practices and security outcomes provide empirical evidence for security frameworks and regulations. Policymakers can leverage these findings to develop guidelines that promote frameworks emphasizing coordinated, measurable practice adoption rather than checklist-based compliance. By aligning regulatory efforts with empirically supported metrics, policy can more effectively incentivize secure development behaviors across the software ecosystem.
Future Work: Our findings highlight the need to better understand the temporal feedback loop between security practices and outcomes. As projects adopt more practices, they may initially report more vulnerabilities and delays in dependency updates and fixing vulnerabilities due to improved detection rather than increased risk. The complexity calls for longitudinal studies that track how individual and aggregated practices affect security outcome metrics. For researchers, future work should model these dynamics in the context of project maturity, contributor activity, and development cycles to clarify whether observed security trends reflect degradation or deliberate, ongoing improvement. A comprehensive understanding of security outcome metrics is an open research challenge; therefore, researchers should explore new security outcome metrics beyond those used in this study.
7.2. Threat to validity
@@
Our regression and causal analyses provide evidence supporting associations and causality between aggregated Scorecard scores, security practices, and security outcomes within our dataset. However, several inherent biases in our dataset must be acknowledged. First, our analyses are based on snapshot data capturing repository Scorecard metric scores and characteristics at a single point in time, lacking temporal granularity. Without longitudinal or time-series analysis, uncertainty remains about whether changes in security practices precede or directly influence observed security outcomes, thereby limiting the generalizability of our findings beyond our dataset. Second, selection bias may be present, as repositories adopting security practices could differ in unmeasured factors such as organizational policies, economic incentives, developer motivations, or project complexity. Although PSM mitigates confounding through observed covariates, unobserved variables may still influence both practices and outcomes.
@@
Apart from that, our dataset exhibits other imbalances, as repositories have security practice scores of either 0, or -1, reflecting the limited adoption of these practices across the npm ecosystem. The real-world imbalance, characterized by the few repositories that adopt security practices, may introduce bias into our models, potentially skewing predictions or causal estimations. However, this imbalance highlights a widespread lack of security practice adoption among npm repositories. Hence, the imbalance reflects the actual state of adoption rather than a methodological flaw, highlighting the need for targeted interventions to increase the adoption of security practices among npm repositories. At the same time, the metric collection also needs improvement, a team may be following a practice, but evidence of that practice may not be captured in the automated metrics.
8. Related Work
This section highlights prior related studies, existing security guidelines, frameworks, and security outcome metrics. Prior studies (zahan2023software, ; sonatype_8, ) have utilized OpenSSF Scorecard metrics alongside vulnerability count data to understand the relationship between security practices and security outcome metrics. Zahan et al. (zahan2023software, ) analyzed 2,422 PyPI and npm packages and found a positive association between Scorecard metrics and vulnerability counts at the package level, but with low explanatory power ($R^{2}$ $<$ 12%). They attributed this to limited dataset size, sparse vulnerability data at the package level, and missing control variables. Sonatype (sonatype_8, ) used a random forest classifier to predict vulnerable repositories based on Scorecard metrics and project attributes. However, binary classification limits understanding of the strength and direction of relationships between practices and outcomes. In contrast, our study employs regression and causal analysis to quantify associations, uses a larger, more representative dataset of 145,817 npm packages, incorporates project-level outcome metrics, and controls for key project characteristics to improve validity and generalizability.
@@ 8. Related Work
Security Practice Frameworks: Executive Order (EO) 14028 emphasized adopting security practices to improve software security. To align with the EO, government agencies and industry experts have introduced various software supply chain security frameworks to enhance software supply chain security. These include the Software Component Verification Standard (SCVS) by OWASP (scvs, ), the Building Security In Maturity Model (BSIMM) (BSIMM, ), the CNCF Technical Advisory Group (TAG) recommendations (CNCF_TAG, ), Supply Chain Levels for Software Artifacts (SLSA) (SLSA, ), NIST Secure Software Development Framework (SSDF) (souppaya2022secure, ), the Open Source Software (OSS) Secure Supply Chain (SSC) Framework by Microsoft (Microsoft_framework, ), NIST’s Cybersecurity Supply Chain Risk Management Practices (NIST800-161, ), the CISA Secure Software Development Attestation Form (CISASecureSoftwareAttestation, ) and OpenSSF Scorecard (Scorecard, ). The Proactive Software Supply Chain Risk Management (P-SSCRM) (Williams2024PSSCRM, ) released in 2024, results from the study of real-world software supply chain risk management initiatives and the union of the 73 security practices in ten government and industry documents (standards and frameworks) mentioned above. In 2025, EO 14144 (EO2025Cybersecurity, ) was issued, emphasizing the need for NIST to evaluate common cyber practices and security outcomes across industry sectors.
@@
Security Outcome Metrics:Various metrics can assess the effectiveness of security practices by tracking security performance over time, generally categorized as threat-oriented, asset/impact-oriented, and vulnerability-oriented (NIST800-55, ; NIST800-30, ). While threat- and asset/impact-oriented metrics offer valuable insights, they are difficult to quantify at scale due to reliance on real-time threat intelligence and complex impact assessments. In contrast, vulnerability-oriented metrics are more accessible and relevant to software security. Among them, severity-based metrics (e.g., CVSS) face challenges such as scoring inconsistencies and reliance on exploitability data (pendleton2016survey, ). Temporal vulnerability metrics, such as vulnerability count, time to detect, time to repair and time to remediate, are widely used to evaluate security (morrison2014mapping, ; sec_outcome, ; CISA-2023, ; morrison2018mapping, ). Morrison et al. (morrison2018mapping, ) conducted a systematic mapping study on software security metrics to evaluate the security properties of software and found that the most cited and most used metric is the vulnerability count (vulnerability metric). Davidson (davidson2009good, ) studied the time to close bug/vulnerability’ (a form of mean time to repair) as a security metric and measured its effectiveness. Beres et al. (beres2009using, ) studied the window of exposure (vulnerability metric), which is the length of the interval of time needed by the security team to remediate a vulnerability once it has been disclosed. However, the scarcity of reported vulnerabilities, compounded by silently fixed issues, limits their reliability in assessing security outcomes (alhazmi2007measuring, ; dunlap2023finding, ). To address these limitations, prior work suggests alternative metrics, such as fault prediction models (shin2013can, ). Additionally, in software reliability, MTTR and MTTU have long been used to measure system resilience and maintenance efficiency (gokhale1999time, ; calvo2023applying, ). While MTTR is often employed in security contexts (MTTR_1, ; MTTR_2, ; MTTR_3, ), its reliance on reported vulnerabilities restricts its applicability. MTTU, by focusing on dependency updates, offers broader coverage but remains constrained by proprietary measurement methods (cogo2019empirical, ).
@@ 9. Conclusion
In this study, we investigated how the aggregate Scorecard score and Scorecard security practices metrics impact security outcome metrics. Our regression model and causal analysis reveal that higher aggregated Scorecard scores are associated with fewer vulnerabilities, shorter MTTR, and shorter MTTU. However, while regression models and causal analysis indicate better outcomes, our segment analysis to understand the impact of project characteristics suggests that factors beyond security practice adoption, such as project maintainability, resource availability, and dependency complexity, influence security outcomes. Feature ranking highlights Code Review, Contributors, License, CI-Tests, and Pinned Dependencies as top security practices associated with security outcomes, though their associations with security outcomes vary in direction. Additionally, MTTU exhibits trends similar to vulnerability count and MTTR, making it a potential proxy metric when vulnerability and MTTR data are unavailable.